umency D Best Bay Credit c8B.A in Risk Man... Degrees&Majors Connect-Classww.wley.m DashboardKMytaisc US Kimmel, Accounting, Ge PRIN. OF ACCT I&II (ACCT 2101& 210 STANDARD view Don Lieberman and concerning the three audit Jobs conducted during November uses job order costing to capture the costs of ts audit jobs. There were no audit jobs in process at the beginning of November. Listed below are data Lynn Brian Mike Direct materials Auditor labor costs Auditor hours 510 $6,000 $280 $7,600$3,575 49 470 91 Overhead costs are applied to jobs on the basis of auditor hours, and the predetermined a overhead rate is $55 per auditor hour. The Lynn job is the only incomplete job at the end of November. Actual overhead for the month was $13,800. (a) Determine the cost of each job. Lynn Brian (b) Indicate the balance of the Service Contracts in Process account at the end of November Balance in service contracts in process accound (c) Calculate the ending balance of the Operating Overhead account for November Balance in operating overtead account Click if you would like toe Show Work for this question: Qeen Show Wenh 6 8
15-12
Answer #1

Answers

  • All working forms part of the answer
  • Requirement ‘a’

Direct Material

Auditor Labor Cost

[Applied] Overhead cost

Total Cost = ANSWER

Lynn

$610

$6,000

$4,070 [74 x $55]

$10,680

Brian

$470

$7,600

$5,005 [91 x $55]

$13,075

Mike

$280

$3,575

$2,695 [49 x $55]

$6,550

  • Requirement ‘b’

Balance in Service Contract in Process account = Cost of unfinished contract = Cost of Lynn Job = $ 10,680

  • Requirement ‘c’

Balance in Operating Overhead account = Difference between Actual Overhead and Applied Overhead

Actual Overhead = $ 13800
Applied Overhead = 4070 + 5005 + 2695 = $ 11,770
Balance = 13800 – 11770 = $ 2,030= ANSWER